OpenStudy (anonymous):

Solve using the substitution method. 18x + 6y = 78 12x + 54y = -48

OpenStudy (anonymous):

first simplify the equations the first is 3x + y =13, the second 2x + 9y = -8. solve the first for y to get y = 13-3x plug y into the second equation to get 2x + 9(13-3x) = -8 2x + 117 -27x = -8 -25x = -185 x = 185/25 == 37/5 == 7.4 then plug x into the equation we solved for y to get y = 13-3(37/5) y = 13 - 111/5 y = -9.2

OpenStudy (anonymous):

I built off of yours so thank you, but I'm not sure how you got the -185? I got -125 so x=-5.

OpenStudy (anonymous):

Final Answer: 18x+6y=78 3x+y=13 12x+54y=-48 2x+9y=-8 y=-3x+13 2x+9y=-8 2x+9(-3x+13)=-8 2x+(-27x)+117=-8 25x=-125 x=-5 3x+y=13 3(-5)+y=13 -15+y=13 y=28 (-5,28)